3x-1/3 - x-3/15 = 2x+3/2
I'm supposed to solve for x, but I just cant seem to figure it out! Help! :) Thank you!!

Answer by lwsshak3(11628)

3x-1/3 - x-3/15 = 2x+3/2
As written, the given equation is made up of 3 separate x-terms and 3 fractions.
Like when solving all equations, place all the x-terms on one side and numbers on the other side
3x-x-2x=3/2+1/3+3/15
combine like terms
0≠3/2+1/3+3/15
Therefore, there is no solution to this equation